A boy is twice as old as his sister, and half as old as their father. In 50 years, his sister will be as old as their father. How old is the boy now?
Answer by lwsshak3(11628)

A boy is twice as old as his sister, and half as old as their father. In 50 years, his sister will be as old as their father. How old is the boy now?
***
let x=boy's age now
x/2=sister's age now
2x=father's age now
..
x/2+50=2x
1.5x=50
x=33-1/3
boy's age now=33-1/3